Initially milligrams of a radioactive . After hours the mass had decayed by the rate of decay is proportional to the amount of substance present at time amount remaining after hours .
1
Expert's answer
2021-07-16T08:01:50-0400

Initially 100 milligrams of a radioactive . After 6 hours the mass had decayed by 3% the rate of decay is proportional to the amount of substance present at time t amount remaining after 24 hours.

The rate of decay is proportional to the amount of substance present at time "t"

"\dfrac{dA}{dt}=kA""\dfrac{dA}{A}=kdt""\int\dfrac{dA}{A}=\int kdt""\ln |A|=kt+\ln C""A=Ce^{kt }"

Initially 100 milligrams of a radioactive



"100=Ce^{k(0)}=>C=100""A(t)=100e^{kt}"


After 6 hours the mass had decayed by 3%



"97=100e^{6k}""6k=\ln0.97""k=\dfrac{1}{6}\ln 0.97"

After 24 hours



"A(24)=100e^{{\ln 0.97 \over 6}(24)}""A(24)=100e^{4\ln 0.97}""A(24)=100(0.97)^4""A=88.529281\ mg"